# Supplementary Information

### v1

@@ Abstract
Single-molecule spectroscopy (SMS) is an exceptionally sensitive technique, but its inherently limited photon budget produces noisy data that can readily lead to subjective analyses, fitting errors, and reduced statistical power, obscuring true subpopulations and their dynamics. Here, we present an unbiased, objective method to cluster two-dimensional single-molecule data and demonstrate it on fluorescence lifetime–intensity correlations. The clustering method is based on Gaussian mixture modeling, with the optimal number of clusters determined through the Bayesian information criterion (BIC). The BIC score per cluster, which displays in general a non-monotonically decreasing trend, presents multiple local minima as candidate solutions for the number of fitted clusters. We also demonstrate the usefulness of statistically grouping resolved levels. The clustering protocol was benchmarked on simulated data and applied to experimental data from the Alexa Fluor 647 dye, QD 605, and the main light-harvesting complexes of plants and cyanobacteria. The combined application of grouping and clustering achieves substantial noise reduction and the identification of relevant, physically meaningful states that would typically be obscured by manual inspection.
@@ Introduction
The pioneering work of Moerner and Kador on single molecules (SMs) (moerner1989optical) paved the way for research into several highly sensitive optical imaging and spectroscopy techniques capable of detecting nanoscale molecular dynamics. Among these are SM Förster resonance energy transfer (smFRET) (roy2008practical, uphoff2010monitoring), time-resolved fluorescence methods that are usually based on time-correlated single photon counting (TCSPC) (gruber2018isolated, bockenhauer2011conformational), single-particle tracking techniques (van2022real, manzo2015review), superresolution localization imaging methods such as Photoactivated Localization Microscopy (PALM)  (hess2006ultra, betzig2006imaging) and Points Accumulation for Imaging in Nanoscale Topography (PAINT) (sharonov2006wide), SM electrophysiology techniques such as single-channel patch-clamp detection (neher1976single, harms2003probing) and nanopore electrophysiology (bayley2001stochastic), and SM force spectroscopy (yen2018improving). SM techniques are uniquely capable of detecting physically relevant subpopulations, which represent distinct functional states that the system can transition between, information that is masked or averaged out in ensemble techniques.
@@
Raw SM experimental data can often be multidimensional depending on the type and purpose of the measurements. Each data point in a multidimensional parameter space could display information such as conformational dynamics, reaction kinetics, diffusion properties, electronic properties, and environmental responses, and these properties are often correlated with one another. The dimensionality of the data usually scales with computational complexity and visualization limitations. As such, subsets comprising only two or three parameters at a time are typically selected. Clustering can be effectively done on these lower-dimensional distributions and serves a dual purpose in denoising the data and isolating different subpopulations, some of which may be statistically rare but interesting. These data clusters may be physically linked to different functional states, for which the switching dynamics between these clusters can easily be determined. Examples include neural spike detection in patch-clamp detection (ghaderi2018electrophysiological), spectral clustering (kruger2017reduced), smFRET common trajectory identification (blanco2015single), and fluorescence intensity–lifetime correlation classification (schlau2015single).
The choice of clustering protocol depends on various factors, including the general observed trend, computational intensiveness, and prior knowledge. Data clustering is generally considered an unsupervised data classification method (although supervised and semi-supervised clustering also exist) whereby an algorithm identifies patterns in the dataset without prior ground-truth labels encoded into the algorithm. Even within this framework, a user can specify conditions under which the algorithm should operate—the maximum number of clusters, for instance—without altering its unsupervised nature. Although the number of clusters may be user-defined, the actual assignment of data points to clusters remains determined entirely by the algorithm, not by any user-supplied ground truth. Subjective clustering, as is the case for K-means and Gaussian clustering (schlau2015single, joshi2021single), may potentially introduce user bias, which could prevent reproducibility and negatively affect biological relevance and accuracy. Computationally intensive clustering, such as Bayesian cluster analysis (see, e.g., Refs. (rubin2015bayesian, griffie2016bayesian)), tends to be slow and may result in overfitting or underfitting the data. Although Refs. (schlau2015single, rubin2015bayesian, griffie2016bayesian) yielded excellent insight into the dynamics of the respective studied systems, the clustering algorithm used either calls for manual selection of the number of clusters, may be computationally expensive, or may overfit the data.
A more robust approach would involve employing a clustering algorithm that is scalable, probabilistic, and free from bias—a requirement rendered even more critical in light of the inherent complexities and limitations of data analysis. SM data involves a considerable amount of inherent noise that arises from sources such as autofluorescence of optical components, scattered light, other background signals, laser fluctuations, detector noise, sample inhomogeneities, and shot noise. This noise can easily shroud weak signals from rare events or transient states. Adding to these uncertainties are possibly non-optimized experimental conditions, which can cause additional experimental error.
@@
Single-molecule spectroscopy (SMS) is an exquisitely sensitive way to detect temporal parameters such as fluorescence brightness, polarization, spectral properties (such as peak position or spectral linewidth), lifetime, and dwell times of a particular state and can be performed on virtually any nanoscale emitter. Examples of such emitters include quantum dots (QDs) (rombach2013blinking, palstra2021intermittency, palstra2021python), fluorescent dyes (zhang2021investigating), and dendrimer-based systems in which fluorophores can be embedded within their highly-branched polymer scaffolds (paulo2010single, kohn2001single). These systems generally exhibit binary switching behavior in their fluorescence intensity and lifetime, characterized by a well-defined bright, long-lifetime state, and a darker, quenched state, often called “on" and “off" states, respectively, a phenomenon known as photoblinking or fluorescence intermittency. However, many systems are known to deviate from such a bimodal behavior. Even single, fluorophore molecules have been shown to display multiple emissive states (wang2013lifetime). For QDs, the number of accessible states depends on their type, structure, and the excitation conditions. For example, type II-IV QDs generally display bimodal lifetime–intensity distributions (palstra2021python, efros2016origin), while perovskite QDs exhibit more continuous distributions (palstra2021intermittency, yarita2017impact). Clustering at the SM level can be a useful tool to decipher and interpret this complex fluorescence behavior (rombach2013blinking, munoz2022memory, pelton2007evidence, chu2024single).
@@
The presence of multiple chromophores in multichromophoric systems contributes significantly to the heterogeneity in their data space. Examples of such systems include the aforementioned dendrimers as well as conjugated polymers, azobenzenes, diarylethenes, synthetic dye aggregates, DsRed, and photonic wires. Photosynthetic light-harvesting complexes form a unique group of multichromophoric systems because of the potential biological significance of data clusters. For example, a subtle change in protein conformation may switch many light-harvesting complexes between a light-harvesting and a photoprotective state (kruger2011conformational), and these transitions have been directly correlated with photoblinking (kruger2012controlled). In addition, these complexes display rich heterogeneity in their intensity–lifetime distributions (schlau2015single) and spectral properties (kruger2017reduced, goldsmith2010watching). It was proposed that the complex relationship between fluorescence intensity and lifetime cannot be adequately studied by traditional simple kinetic models (goldsmith2010watching). The heterogeneity and dynamics of light-harvesting complexes make them an excellent testbed for SMS clustering algorithms.
@@
Typically, a probability distribution is used to visually reveal relevant states of an isolated SM system (schlau2015single, santoso2010characterizing, kalinin2008characterizing). These distributions can then be clustered based on the number of observed dense regions using clustering algorithms such as Gaussian mixture modeling (GMM) (mattana2022automated, rubbens2021phenogmm) or K-means clustering (wang2013lifetime), although any clustering algorithm could be used (xu2005survey, hartigan1975clustering). Each resulting cluster provides insight into the dynamics of the distinct fluorescence states and may enable inference of information such as the dwell times and switching rates between different states. There are many methods to determine these switching rates, such as the maximum-likelihood estimation of observables (schmid2016single) or inverse mean dwell time (schuler2013single), though the empirical definition of the switching rate is most universal (chung2018transition).
@@
We have developed a general-purpose, open-source Python-based software application for SMS data analysis, called Full SMS (botha2024advanced), which performs, i.a., analysis of intensity time traces using a statistically robust change-point analysis (watkins2005detection). The software also allows statistical grouping of the resolved intensity levels through a combination of agglomerative hierarchical clustering and expectation maximization clustering (watkins2005detection), which significantly improves the fitting of intensity levels and lifetimes. The grouping, which also features a type of clustering, is technically capable of identifying physical states, although it may not be statistically plausible to use only grouping as an objective means to identify these states in a lifetime–intensity distribution. A more rigorous and versatile analysis method is required.
In this work, we present an extension of the former by introducing a robust, multistate objective clustering protocol on experimental post-processed data from the Full SMS software (botha2024advanced). Similar to Full SMS, the repository is publicly hosted on GitHub and can be accessed at https://github.com/BioPhysicsUP/Clustering-Protocol. To demonstrate the power of this clustering algorithm, we tested it on simulated data for benchmarking, followed by the following case studies: randomly oriented fluorescent dye molecules (Alexa Fluor 647) exhibiting rapid photoblinking, QD 605 data, which potentially also display multistate behavior (munoz2022memory), and two light-harvesting complexes as examples of multichromophoric systems.
@@
The experimental setup is described in Ref. (kyeyune2019strong), with some minor modifications. Briefly, a ps pulsed supercontinuum laser (Fianium, SC400-4-PP), with a repetition rate of $40$ MHz, was used as the illumination source, where the central wavelength of excitation was selected using an acousto-optical tunable filter (Crystal Technology). The excitation beam was circularly polarized before passing through a spatial filter to isolate the $\text{TEM}_{00}$ mode, and thereafter focused into a near-diffraction-limited spot by a 1.45-NA Nikon oil-immersion objective in an epifluorescent configuration. The fluorescence photons were captured by the same objective, focused through a confocal pinhole using an appropriately selected dichroic beamsplitter, spectrally cleaned using a suitable long-pass filter, and detected by a single-photon avalanche photodiode (PD-050-CTE, Micro Photon Devices, Bolzano, Italy, IRF $\sim 128$ ps).
@@
Simulated datasets were generated to benchmark the clustering algorithm, the code of which is available at https://github.com/BioPhysicsUP/SMS-Simulations. The inherent noise in SMS experiments includes the fundamental shot noise (obeying Poisson statistics), detector noise (obeying Gaussian statistics), and ambient background noise. The latter arises from stray light sources such as scattering and autofluorescence from optical components and appears as a constant background level, where each stray light source is also affected by shot noise. Background noise is especially problematic for the lowest-intensity populations, where it significantly reduces the signal-to-noise ratio. The detector noise includes dark counts and readout noise, but can be countered by taking a background measurement and subtracting that level from the intensity data. The shot noise is inherent due to the discrete nature of photons arriving at random times. This is generally the major noise contributor in experiments and cannot be subtracted. To simulate the data, random lifetime–intensity populations were generated by defining a sampling population with a user-specified mean and standard deviation (SD) for the lifetime and intensity of each state. The SD was set to $\leq 10\%$ of the corresponding mean. Dwell times were then generated on a per-particle basis: for unquenched (“on") states, they were drawn from a truncated power-law distribution (see Eq. (1)) to model heavy-tailed events arising from trap states, whereas for quenched (“off") states, a simple power-law distribution was used.
The truncated power-law function is given by
@@
where $\alpha$ is the power-law constant. The truncated power law is characterized by a high probability of short dwell-time events and a low probability of long dwell times, with $\tau_{c}$ acting as the exponential cutoff time after which the exponential decay dominates. For this particular study, we set $\alpha=1.2$ and $\tau_{c}=20$ s, and for each particle, dwell times were drawn from this distribution, where the corresponding intensities and lifetimes were sampled from the defined means and SDs, thereby simulating states with normally distributed noise. Finally, Poisson noise (representing shot noise) was added to the intensities to yield simulated two-state intensity traces.
@@
The experimental data were recorded in time-tagged time-resolved (TTTR) mode, stored in HDF5 format, and analyzed using Full SMS (botha2024advanced). Data screening was the first step, whereby particles bleached (visible as an irreversible loss of fluorescence) during the raster scan were omitted, photon bursts were removed, and only the brightest particles were selected. The time traces of the remaining particles were then analyzed within a specific time region until partial or complete photodegradation was observed. The fluorescence intensity levels were then resolved and subsequently grouped using the software’s native clustering functionality, which provided the statistically most probable levels. The fluorescence lifetime of each grouped intensity level was determined using $\chi^{2}$ and Durbin-Watson goodness-of-fit (durbin1950testing) parameters.
The grouped lifetime–intensity data output of the Full SMS software was further analyzed in a raw Python environment to perform additional data clustering. This clustering algorithm (hereafter referred to as the clustering protocol) objectively determines the optimal number of identifiable populations (states) to which a single particle (object) can switch. Data were clustered using GMM, assuming that all points in the distribution were drawn from a mixture of Gaussian distributions with unknown parameters. For each GMM cluster, a mean vector and a covariance matrix were estimated, representing the cluster center and the shape/spread of the data. Data points were assigned according to the highest likelihood of belonging to a cluster. The optimal number of GMM clusters was determined using the Bayesian information criterion (BIC), a statistical measure that balances model complexity with model fit (bouman1997cluster, neath2012bayesian). The BIC value is given by
@@
where $k$ represents the number of parameters, $n$ the number of data points, and $L$ the likelihood fitness of the model. The BIC value was determined for each number of GMM clusters, and the best fit corresponded to the lowest BIC value. Ellipses are drawn around each cluster center, each representing confidence contours for each fitted GMM component. For a two-dimensional Gaussian with mean $\vec{\mu}$ and positive semi-definite covariance matrix $\Sigma$, the contour is defined by the squared Mahalanobis distance:
$(\vec{x}-\vec{\mu})^{T}\Sigma^{-1}(\vec{x}-\vec{\mu})=\chi_{2}^{2}({p}),$ (4)
@@
where $\vec{x}$ is a point on the lifetime–intensity plot, and $\chi^{2}_{2}(p)$ denotes the quantile of the chi-squared distribution with two degrees of freedom at probability $p$. This ellipse encloses a fraction $p$ of the probability mass around the cluster center of each GMM component. Throughout this study, the confidence level $p$ was selected empirically to minimize cluster overlap while preserving adequate information about cluster spread. Once the optimal number of populations was determined, other relevant information, such as the switching rates and switching frequencies between states, was extracted.

### v2

@@ Abstract
Single-molecule spectroscopy (SMS) is an exceptionally sensitive technique, but its inherently limited photon budget produces noisy data that can readily lead to subjective analyses, fitting errors, and reduced statistical power, obscuring true subpopulations and their dynamics. Here, we present an unbiased, objective method to cluster two-dimensional single-molecule data and demonstrate it on fluorescence lifetime–intensity correlations. The clustering method is based on Gaussian mixture modeling, with the optimal number of clusters determined through information criteria (the Akaike and Bayesian information criteria and integrated completed likelihood) and supplemented by cluster quality metrics such as average cluster tightness and the fraction of points outside confidence ellipses, which guide the selection of statistically robust and physically meaningful clusters. The protocol was benchmarked on simulated datasets spanning clean, smeared, and noisy overlap-limited regimes, and applied to experimental data from Alexa Fluor 647 and QD 605. This approach reliably recovers relevant subpopulations even in the presence of noise and overlapping distributions, providing an objective framework for analyzing single-molecule heterogeneity, with limitations arising primarily under severe geometric overlap or extreme state-occupancy imbalance where distinct populations are no longer separable.
@@ Introduction
The pioneering work of Moerner and Kador on single molecules (SMs) (1) paved the way for research into several highly sensitive optical imaging and spectroscopy techniques capable of detecting nanoscale molecular dynamics. Among these are SM Förster resonance energy transfer (smFRET) (2, 3), time-resolved fluorescence methods that are usually based on time-correlated single photon counting (TCSPC) (4, 5), single-particle tracking techniques (6, 7), superresolution localization imaging methods such as Photoactivated Localization Microscopy (PALM)  (8, 9) and Points Accumulation for Imaging in Nanoscale Topography (PAINT) (10), SM electrophysiology techniques such as single-channel patch-clamp detection (11, 12) and nanopore electrophysiology (13), and SM force spectroscopy (14). SM techniques are uniquely capable of detecting physically relevant subpopulations, which represent distinct functional states that the system can transition between, information that is masked or averaged out in ensemble techniques.
@@
Raw SM experimental data can often be multidimensional depending on the type and purpose of the measurements. Each data point in a multidimensional parameter space could display information such as conformational dynamics, reaction kinetics, diffusion properties, electronic properties, and environmental responses, and these properties are often correlated with one another. The dimensionality of the data usually scales with computational complexity and visualization limitations. As such, subsets comprising only two or three parameters at a time are typically selected. Clustering can be effectively done on these lower-dimensional distributions and serves a dual purpose in denoising the data and isolating different subpopulations, some of which may be statistically rare but interesting. These data clusters may be physically linked to different functional states, for which the switching dynamics between these clusters can easily be determined. Examples include neural spike detection in patch-clamp detection (15), spectral clustering (16), smFRET common trajectory identification (17), and fluorescence lifetime–intensity correlation classification (18).
Clustering is a central tool in SMS (19, 20, 21, 22) for identifying recurring photophysical populations in heterogeneous datasets, yet most commonly used approaches require a priori assumptions that limit objectivity and reproducibility. Widely used methods, such as K-means and Gaussian clustering (23, 18, 24), are nominally unsupervised but remain subjective in practice, as the user must specify key parameters—most notably the desired number of clusters—which can strongly bias the outcome and interpretation. K-means further imposes restrictive assumptions of spherical clusters with equal variance and no covariance structure, conditions that are rarely met in SMS data characterized by heterogeneous noise sources, anisotropic correlations, and limited photon statistics. In addition, heuristic model-selection approaches commonly used with K-means, such as the elbow method based on the within-cluster sum of squares (WCSS) (25), rely on objective functions that decrease monotonically with increasing cluster number, rendering elbow identification highly subjective and unreliable in ambiguous or noisy regimes typical of SMS. Gaussian mixture models (GMMs) (26, 27) extend Gaussian clustering by modeling the data as a probabilistic mixture of multivariate Gaussian components rather than assigning hard labels, allowing cluster shape, orientation, and covariance to be explicitly captured. Importantly, GMMs are inherently compatible with information-criterion (IC) frameworks such as the Akaike information criterion (AIC) (28, 29), Bayesian information criterion (BIC) (30, 31), and integrated completed likelihood (ICL) (32), which enables near-automated, statistically grounded selection of the optimal number of components. Fully Bayesian clustering (33, 34) approaches, including Dirichlet Process GMMs (DPGMMs, based on Bayesian inference—see, e.g., Ref. (35)) in principle eliminate the need for manual model selection by treating the number of clusters as a random variable. In practice, however, these methods are computationally expensive, sensitive to prior specification, and prone to over-partitioning noisy, low-occupancy regions. These limitations are particularly problematic for photon-limited SMS data, where diffuse background fluctuations can be spuriously promoted to physically meaningful states. Although Bayesian covariance estimation can be more stable for small samples, it is inherently biased toward the prior and may generate numerous micro-clusters, undermining interpretability and reproducibility.
@@
SMS is an exquisitely sensitive way to detect temporal parameters such as fluorescence brightness, polarization, spectral properties (such as peak position or spectral linewidth), lifetime, and dwell times of a particular state, and can be performed on virtually any nanoscale emitter. Examples of such emitters include quantum dots (QDs) (36, 37, 38), fluorescent dyes (39), and dendrimer-based systems in which fluorophores can be embedded within their highly-branched polymer scaffolds (40, 41). These systems generally exhibit binary switching behavior in their fluorescence intensity and lifetime, characterized by a well-defined bright, long-lifetime state, and a darker, quenched state, often called “on" and “off" states, respectively, a phenomenon known as photoblinking or fluorescence intermittency. However, many systems are known to deviate from such a bimodal behavior. Even single, fluorophore molecules have been shown to display multiple emissive states (23). For QDs, the number of accessible states depends on their type, structure, and the excitation conditions. For example, type II-IV QDs generally display bimodal lifetime–intensity distributions (38, 42), while perovskite QDs exhibit more continuous distributions (37, 43). Clustering at the SM level can be a useful tool to decipher and interpret this complex fluorescence behavior (36, 44, 45, 46). At the level of clustered observables, locally dense populations are well approximated by Gaussian distributions, a reasonable assumption in SMS where multiple independent noise contributions lead to approximately Gaussian local behavior via the central limit theorem.
@@
The presence of multiple chromophores in multichromophoric systems contributes significantly to the heterogeneity in their data space. Examples of such systems include the aforementioned dendrimers as well as conjugated polymers, azobenzenes, diarylethenes, synthetic dye aggregates, DsRed, and photonic wires. Photosynthetic light-harvesting complexes form a unique group of multichromophoric systems because of the potential biological significance of data clusters. For example, a subtle change in protein conformation may switch many light-harvesting complexes between a light-harvesting and a photoprotective state (47), and these transitions have been directly correlated with photoblinking (48). In addition, these complexes display rich heterogeneity in their lifetime–intensity distributions (18) and spectral properties (16, 49). For example, allophycocyanin, a core subunit of the main light-harvesting complex in cyanobacteria and red algae, has been shown to exhibit six distinct emissive states and a complex relationship between their fluorescence intensity and lifetime that precludes the use of simple, traditional kinetic models (49).
@@
Typically, a probability distribution is used to visually reveal relevant states of an isolated SM system (18, 50, 51). These distributions can then be clustered based on the number of observed dense regions using clustering algorithms such as GMM. Each resulting cluster provides insight into the dynamics of the distinct fluorescence states and may enable inference of information such as the dwell times and switching rates between different states. There are many methods to determine these switching rates, such as the maximum-likelihood estimation of observables (52) or inverse mean dwell time (53), though the empirical definition of the switching rate is most universal (54).
@@
We have developed a general-purpose, open-source Python-based software application for SMS data analysis, called Full SMS (55), which performs, i.a., analysis of intensity time traces using a statistically robust change-point analysis (56). The software also allows statistical grouping of the resolved intensity levels through a combination of agglomerative hierarchical clustering and expectation maximization clustering (56), which significantly improves the fitting of intensity levels and lifetimes. While this grouping can identify physical states, it may not be statistically sufficient as an objective means to define states in a lifetime–intensity distribution, highlighting the need for a more versatile analysis method.
In this work, we introduce an objective, multistate clustering protocol for SMS data and demonstrate it on two-dimensional fluorescence lifetime–intensity distributions, with the lifetime and intensity data output by Full SMS either before or after intensity-level grouping. While we here employ the Parquet file format for efficient storage and reproducible data handling, the clustering method is software-agnostic and can be applied to any SMS dataset conforming to the required structure. The protocol also supports export in multiple formats. Whereas the earlier work  (55) used statistical grouping primarily descriptively, the present study formalizes clustering as a statistically guided inference problem by systematically combining GMMs with multiple IC and auxiliary quality metrics, enabling objective identification of dominant photophysical populations across complex datasets. To assess robustness and practical limits, we perform controlled simulations spanning different state constructions, blinking models, and noise contributions, followed by application to two experimental datasets, viz., a rapidly blinking, randomly oriented Alexa Fluor 647 molecules dataset, as well as a small-sized QD 605 dataset (44). The clustering protocol is publicly available via GitHub at https://github.com/BioPhysicsUP/Clustering-Protocol.
@@
The experimental setup is described in Ref. (57), with some minor modifications. Briefly, a ps pulsed supercontinuum laser (Fianium, SC400-4-PP), with a repetition rate of $40$ MHz, was used as the illumination source, where the central wavelength of excitation was selected using an acousto-optical tunable filter (Crystal Technology). The excitation beam was circularly polarized before passing through a spatial filter to isolate the $\text{TEM}_{00}$ mode, and thereafter focused into a near-diffraction-limited spot by a 1.45-NA Nikon oil-immersion objective in an epifluorescent configuration. The fluorescence photons were captured by the same objective, focused through a confocal pinhole using an appropriately selected dichroic beamsplitter, spectrally cleaned using a suitable long-pass filter, and detected by a single-photon avalanche photodiode (PD-050-CTE, Micro Photon Devices, Bolzano, Italy, IRF $\sim 128$ ps).
@@
Simulated datasets were generated to benchmark the clustering algorithm, the code of which is available at https://github.com/BioPhysicsUP/SMS-Simulations. The inherent noise in SMS experiments includes the fundamental shot noise (obeying Poisson statistics), detector noise (obeying Gaussian statistics), and ambient background noise. The latter arises from stray light sources such as scattering and autofluorescence from optical components and appears as a constant background level, where each stray light source is also affected by shot noise. Background noise is especially problematic for the lowest-intensity populations, where it significantly reduces the signal-to-noise ratio. The detector noise includes dark counts and readout noise, but can be accounted for by taking a background measurement and subtracting that level from the intensity data. The shot noise is inherent due to the discrete nature of photons arriving at random times. This is generally the major noise contributor in experiments and cannot be subtracted. To simulate the data, random lifetime–intensity populations were generated by defining a sampling population with a user-specified mean and standard deviation (SD) for the lifetime and intensity of each state. Dwell times were then generated on a per-particle basis. In the main text, both quenched (“off”) and unquenched (“on”) states were drawn from exponential distributions (see Eq. (1)). In the SI, we explored more complex scenarios in which the “off” dwell times follow a power-law distribution, while the “on” dwell times are drawn from either an exponential, a simple power-law (Eq. (2)), or a truncated power-law (Eq. (3)). This demonstrates the capability of the clustering protocol to handle datasets governed by different underlying statistical models.
The exponential dwell-time probability distribution function is given by
$P(t)=\frac{1}{\tau}e^{-t/\tau},$ (1)
where $t$ is the dwell time and $\tau$ is a sampled mean dwell time, set to $\tau=2$ s for the off-state and $\tau=4$ s for the on-states in this study. The simple power-law probability distribution is given by
$P(t)=t^{-\alpha},$ (2)
where $\alpha$ is the power-law constant, set to $\alpha=1.2$ for this study. The truncated power-law function is given by
@@
where $\tau_{c}$ acts as the exponential cutoff time after which the exponential decay starts to dominate. The truncated power law is characterized by a high probability of short dwell-time events and a low probability of long dwell times. For this particular study, we set $\alpha=1.2$ and $\tau_{c}=20$ s.
For each particle, dwell times were drawn from these distributions, where the corresponding intensities and lifetimes were sampled from the defined means and SDs, thereby simulating states with normally distributed noise. Finally, Poisson noise (representing shot noise) was added to the intensities to yield the desired simulated intensity traces.
@@
The experimental data were recorded in time-tagged time-resolved (TTTR) mode, stored in HDF5 format, and analyzed using Full SMS (55). Data screening was the first step, whereby particles bleached (visible as an irreversible loss of fluorescence) during the raster scan were omitted, photon bursts were removed, and only the brightest particles were selected. The time traces of the remaining particles were then analyzed within a specific time region until partial or complete photodegradation was observed. The fluorescence intensity levels were then resolved and subsequently grouped using the software’s native clustering functionality, which provided the statistically most probable levels. The fluorescence lifetime of each grouped intensity level was determined using $\chi^{2}$ and Durbin-Watson (58) goodness-of-fit parameters.
The correlated lifetime–intensity data output of the Full SMS software was further analyzed in a raw Python environment to perform additional data clustering. This clustering algorithm (hereafter referred to as the clustering protocol) objectively determines the optimal number of identifiable populations (states) to which a single particle (object) can switch. Data were clustered using GMM, assuming that all points in the distribution were drawn from a mixture of Gaussian distributions with unknown parameters. For each GMM cluster, a mean vector and a covariance matrix were estimated, representing the cluster center and the shape/spread of the data. Data points were assigned according to the highest likelihood of belonging to a cluster. The optimal number of GMM clusters was determined using three IC metrics—AIC, BIC, and ICL—supplemented by cluster-quality measures including relative IC gain, average cluster tightness, and the fraction of points outside confidence regions. This combination enables reproducible and minimally biased selection of the number of clusters.
The AIC score was calculated as
$\text{AIC}=2k-2\cdot\text{ln}(L),$ (5)
where $k$ is the number of free parameters and $L$ the maximized likelihood. AIC penalizes model complexity weakly and therefore tends to favor models with a larger number of clusters, emphasizing predictive accuracy. The BIC score, in comparison, is defined as
@@
where $n$ is the number of data points. BIC penalizes model complexity more strongly than AIC, favoring parsimonious solutions consistent with Occam’s razor. ICL extends the BIC formulation by explicitly incorporating cluster assignment certainty:
$\text{ICL}=\text{BIC}-\sum_{i=1}^{n}\sum_{j=1}^{k}\tau_{ij}\cdot\ln(\tau_{ij}),$ (7)
where $\tau_{ij}$ denotes the posterior probability that data point $i$ belongs to cluster $j$. The additional entropy term penalizes overlapping or uncertain assignments, favoring well-separated clusters. Although BIC is most commonly used in SMS studies, including AIC and ICL provides complementary information on predictive accuracy, model parsimony, and cluster separation. All three IC metrics are derived from the GMM likelihood framework and therefore integrate naturally with probabilistic cluster modeling.
IC scores were computed for each candidate number of GMM clusters. The conventional approach is to identify the optimal solution as the model yielding a global minimum IC value. However, when these flexible mixture models are applied to noisy SM data, which is subject to heterogeneity, the IC values frequently decrease with increasing cluster number, and reach a global minimum at unrealistically large numbers of components, leading to possible overpartitioning and substantial computational cost, while offering little physical interpretability (as illustrated in the experimental results discussed in this study). To mitigate this, our algorithm focuses on substantial improvements in model fit as quantified by decreases in the IC, rather than on absolute IC minimization. Specifically, we computed the relative change in the IC score between successive cluster numbers, normalized to the maximum observed change. Here, an IC gain is defined as a pronounced reduction in the IC value upon increasing the number of clusters, yielding an elbow point, which denotes a point at which a steeper gradient transitions to a markedly shallower gradient. Cluster numbers producing sharp IC decreases, whose gradients exceed a threshold set by one standard deviation above the mean IC gradient, were identified as prominent local elbow points; once the IC gradient falls below this threshold, that cluster is no longer considered to provide statistically meaningful improvement. These prominent local elbow points were flagged as candidate solutions, generating a set of statistically grounded cluster choices. Importantly, experimental datasets may exhibit multiple such prominent elbows, reflecting competing yet plausible descriptions of the underlying state structure. Consequently, for such datasets, IC analysis alone is insufficient to select a unique optimal solution. From the resulting candidate set, final model selection was therefore performed using complementary cluster-quality metrics (vide infra), which quantify cluster compactness, separation, and physical plausibility. This two-stage procedure prioritizes models that achieve both statistically meaningful IC gains and physically interpretable clustering, avoids reliance on subjective visual inspection, and mitigates systematic overfitting while remaining sensitive to genuine molecular state structure.
@@
Ellipses were drawn around each cluster center to represent confidence contours for each fitted GMM component. For a two-dimensional Gaussian with mean $\vec{\mu}$ and positive semi-definite covariance matrix $\Sigma$, the contour is defined by the squared Mahalanobis distance:
$D^{2}=(\vec{x}-\vec{\mu})^{T}\Sigma^{-1}(\vec{x}-\vec{\mu})=\chi_{2}^{2}({p}),$ (8)
@@
where $\vec{x}$ is a point on the lifetime–intensity plot, and $\chi^{2}_{2}(p)$ denotes the quantile of the chi-squared distribution with two degrees of freedom corresponding to the confidence level $p$. This ellipse encloses a fraction $p$ of the probability mass around the cluster center of each GMM component.
These confidence ellipses also serve as quantitative guides in cluster evaluation. Cluster tightness was defined as the mean squared Mahalanobis distance of points to their assigned cluster centers, averaged across all clusters ($\langle D^{2}\rangle$). For an ideal two-dimensional Gaussian, the expected value is approximately equal to the dimensionality, i.e., $\langle D^{2}\rangle\approx 2$. Deviations indicate overly compact (<2<2) or loose (>2>2) clusters. The fraction of points outside the confidence ellipse (FO) was also computed, with deviations from the expected value, 1p1-p, indicating overlap, poor separation, or model mismatch. Together, these metrics provide intuitive checks on clustering robustness. Throughout this study, $p$ was selected empirically to minimize cluster overlap while preserving adequate information about cluster spread. Candidate clustering solutions were selected based on a balance between statistical parsimony, cluster tightness, and information gain across multiple metrics, minimizing subjective bias in cluster number determination.
Although many other clustering algorithms could be employed in principle (59, 60), careful choice of an appropriate algorithm is key. The specific data structure is an important consideration. For lifetime–intensity data, the point density varies strongly with intensity, lifetime, and photon count, causing some physically meaningful low-intensity or short-lifetime states to be misclassified as noise or fragmented across clusters in clustering paradigms such as DBSCAN. More generally, SMS observables are naturally described by elliptical distributions with non-zero covariance, rendering GMMs an appropriate clustering choice as they accommodate this structure through full covariance estimation and probabilistic assignment. We therefore used GMM as the primary clustering framework in this work, while comparing the performance of the DPGMM and K-means clustering methods on selected datasets.
In Bayesian formulations of GMMs such as DPGMMs, prior distributions are placed on the mixture weights and component parameters, allowing the effective number of clusters to be inferred by suppressing components with negligible posterior weight. In this study, DPGMM was implemented by specifying an initial upper bound on the number of mixture components, chosen deliberately to be slightly higher than the expected number of states, after which the model redistributes weight among clusters during inference. This reflects a realistic experimental scenario in which the true number of populations is unknown, and a robust clustering algorithm should collapse redundant components. The resulting solution depends sensitively on the choice of hyperparameters, particularly the concentration parameter, and under inappropriate prior specification, DPGMMs may assign low-occupancy or spurious components (called “ghost” clusters). These solutions were therefore evaluated with respect to both the inferred number of clusters and the emergence of such ghost components.
By contrast, K-means clustering was applied as a distance-based reference method. The algorithm performs hard assignments by minimizing the squared Euclidean distance (i.e., WCSS) between data points and cluster centroids, assigning each point $\vec{x}_{i}$ to the cluster $k$, with cluster center $\vec{\mu}_{k}$, such that $|\vec{x}_{i}-\vec{\mu}_{k}|^{2}$ is minimized. WCSS values were computed for cluster numbers ranging from one up to a predefined maximum and plotted as an elbow-point curve. The optimal number of clusters was identified as the point at which the WCSS trend transitions from a steep to a shallow gradient, determined by fitting linear trends to the low- and high-cluster-number regimes and locating their intersection. Cluster boundaries are therefore defined by loci of equal distance between centroids, resulting in linear decision surfaces corresponding to Voronoi partitions of the feature space. This formulation assumes spherical clusters with equal variance and neglects covariance between observables, properties that are generally incompatible with lifetime–intensity data exhibiting correlated noise and unequal spread. Under these conditions, distance-based hard clustering is expected to yield unstable or biased partitions relative to covariance-aware probabilistic models.
Once the optimal number of populations was determined, clustering performance was evaluated using the Adjusted Rand Index (ARI) (61), which quantifies agreement between the true simulated state labels and the cluster assignments returned by the GMM at the level of individual dwell segments. Unlike metrics based solely on cluster-center matching, ARI directly measures point-level assignment accuracy and explicitly penalizes both merged and missing clusters, making it well-suited for situations in which the recovered number of clusters differs from the true number of physical states. An ARI value of unity indicates perfect recovery, while values approaching zero correspond to random assignment. Other relevant information, such as the switching rates and switching frequencies between states, could then be extracted.